# A Swift and Omnidirectional Formation Approach based on Hierarchical Reorganization

### v1

@@ IV-C Power-Centric Topology Switching
The control strategy outlined in Section IV-B reveals a temporal difference in the expected position updates across various hierarchical formation levels. The follower’s position, inherently dependent on topological relationships, exhibits a coupling effect between the formation’s inherent topological structures and the temporal difference, affecting the control accuracy. Considering the dynamic change of individual roles within the hierarchical reorganization, invariant topological connections can compromise control accuracy. In response, this section delves into a power-centric topological switching approach designed to mitigate the influence of topological connections in tandem with the hierarchical reorganization process.
@@
(Generic Graph Configuration [33]) For the dynamic formation $\mathcal{G}(\mathcal{V}(t),\mathcal{E}(t))$, if and only if $\mathcal{G}(\mathcal{V}(t),\mathcal{E}(t))$ is $(d+1)-rooted$, the stress matrix $\Omega$ is semi-positive and $rank(\Omega)=n-d-1$.
@@
Theorem IV.4.
(Topology Configuration for Omnidirectional Affine Localizability) Under the Assumption 4, for the dynamic formation $\mathcal{G}(\mathcal{V}(t),\mathcal{E}(t))$, when each follower is topologically connected only to leaders, $\mathcal{G}(\mathcal{V}(t),\mathcal{E}(t))$ can achieve omnidirectional affine localizability.
@@
$\small\boldsymbol{D}_{fl}=\begin{bmatrix}-\mathbf{1}\in\mathbf{R}^{1\times n_{ef}^{1}}&\mathbf{0}\in\mathbf{R}^{1\times n_{ef}^{2}}&...&\mathbf{0}\in\mathbf{R}^{1\times n_{ef}^{n_{f}}}\\ \mathbf{0}\in\mathbf{R}^{1\times n_{ef}^{1}}&-\mathbf{1}\in\mathbf{R}^{1\times n_{ef}^{2}}&...&\mathbf{0}\in\mathbf{R}^{1\times n_{ef}^{n_{f}}}\\ ...&...&...&...\\ \mathbf{0}\in\mathbf{R}^{1\times n_{ef}^{1}}&\mathbf{0}\in\mathbf{R}^{1\times n_{ef}^{2}}&...&-\mathbf{1}\in\mathbf{R}^{1\times n_{ef}^{n_{f}}}\end{bmatrix}$
@@
Denote $\boldsymbol{\bar{r}}=\begin{bmatrix}\boldsymbol{r}_{1}^{T}&\boldsymbol{r}_{2}^{T}&...&\boldsymbol{r}_{n_{l}}^{T}\\ \mathbf{1}&\mathbf{1}&...&\mathbf{1}\end{bmatrix}^{T}$, then $\boldsymbol{\bar{r}}^{T}\boldsymbol{\omega}=0$. However, according to the Lemma 1 in [8], $rank(\boldsymbol{\bar{r}})=d+1$, which implies $\boldsymbol{\bar{r}}^{T}$ is of row-full rank. Then, from the properties of vector spaces, only zero vectors exist in the zero space of $\boldsymbol{\bar{r}}^{T}$. According to the definition of $\omega$ in Eq.(2), $\sum_{j=1}^{n_{ef}^{i}}|\omega_{i}^{j}|>0$, which contradicts the above inference obviously. ∎
According to Theorem IV.4, when the hierarchical reorganization is completed through Section IV-A, the topological connection structure in the current configuration is reset, such that each follower establishes a topological connection only with leaders. An example of topological switch is shown in Fig. 5.
@@ V-A Simulation
Table I illustrates the impact of hierarchical reorganizations on the different drone path lengths. When employing the formation strategies, Drone 2 exhibits the longest motion path, measuring $85.33m$, while Drone 4 covers the shortest distance at $81.22m$, resulting in a difference of $4.11m$. Conversely, without hierarchical reorganizations, Drone 1 records the longest motion path within the formation at $90.30m$, while Drone 3 covers the shortest distance at $72.79m$, yielding a substantial difference of $18.01m$. This signifies that the omnidirectional affine formation strategies can effectively equalize the motion distances of drones, thereby significantly enhancing both the coordination within the formation and the formation’s overall endurance.
@@
With hierarchical reorganization 82.38 84.86 85.33 82.73 81.22
W/O hierarchical reorganization[8] 88.95 90.30 80.38 72.29 79.56
@@
With hierarchical reorganization 2.12 2.18 2.20 2.12 2.07
W/O hierarchical reorganization[8] 2.39 2.42 2.14 1.94 2.14
Figure 7: Environment setup.
Figure 8: The experimental process of formation flight with patterns ’1’ and ’8’.
Table II presents the influence of hierarchical reorganizations on the average motion speeds of different drones. When employing the omnidirectional affine formation strategy, Drone 2 achieves the highest motion speed, recording a speed of $2.20m/s$, while Drone 4 exhibits the slowest motion speed at $2.07m/s$, resulting in a difference of $0.13m/s$. In contrast, without the hierarchical reorganizations, Drone 1 attains the highest motion speed within the cluster at $2.42m/s$, while Drone 3 moves at the slowest speed of $1.94m/s$, resulting in a more substantial difference of $0.48m/s$. Furthermore, the overall motion durations for both approaches are $38.93s$ and $37.31s$, respectively. This indicates that, with similar overall motion durations, the omnidirectional affine formation strategy significantly reduces the average speed disparities among individual entities, allowing the swarm to maintain similar motion speeds. Consequently, this enhances both the in-team coordination and flexibility of the formation.
TABLE III: The average velocity, path length and times of serving as leaders for different drone in Exp. 1 and Exp. 2
Drone 0 Drone 1 Drone 2 Drone 3 Drone 4
Exp. 1 ${}^{0}\bar{v}(m/s)$ ${}^{1}\bar{v}(m/s)$ ${}^{2}\bar{v}(m/s)$ ${}^{3}\bar{v}(m/s)$ ${}^{4}\bar{v}(m/s)$
0.62 0.63 0.59 0.60 0.60
${}^{0}path(m)$ ${}^{1}path(m)$ ${}^{2}path(m)$ ${}^{3}path(m)$ ${}^{4}path(m)$
14.59 14.73 13.34 14.38 14.73
${}^{0}Times$ ${}^{1}Times$ ${}^{2}Times$ ${}^{3}Times$ ${}^{4}Times$
2/3 2/3 3/3 2/3 3/3
Exp. 2 ${}^{0}\bar{v}(m/s)$ ${}^{1}\bar{v}(m/s)$ ${}^{2}\bar{v}(m/s)$ ${}^{3}\bar{v}(m/s)$ ${}^{4}\bar{v}(m/s)$
0.62 0.59 0.56 0.59 0.62
${}^{0}path(m)$ ${}^{1}path(m)$ ${}^{2}path(m)$ ${}^{3}path(m)$ ${}^{4}path(m)$
15.70 15.09 14.20 15.26 15.77
${}^{0}Times$ ${}^{1}Times$ ${}^{2}Times$ ${}^{3}Times$ ${}^{4}Times$
3/3 2/3 3/3 3/3 2/3
@@ V-B Experimental Setup
To verify the performance of the proposed algorithm, we conduct several indoor experiments, as shown in Fig. 7. Specifically, Fig. 7(a) depicts a scenario set within an obstacle-free dark environment, where five drones execute formation flights in the ${}^{\prime}1^{\prime}$ and ${}^{\prime}8^{\prime}$ patterns. Fig. 7(b) and Fig. 7(c) present environments featuring a hoop obstacle and a natural obstacle, respectively. In Fig. 7(b), the quintet of drones performs a coordinated obstacle avoidance flight through the hoop, whereas in Fig. 7(c), they engage in formation flights with free transformations within the natural obstacle environment. The dimensions (length×$\times$×width×$\times$×height) of each drone are $23cm\times 23cm\times 14cm$. The IMU is embedded in each drone’s flight control hardware, referred to as Pixhawk®. Moreover, the NOKOV® motion capture system provides the global position measurements, and an UP Board 4000® computing board running ROS is adopted as the onboard computer. A small TP-LINK® router model TL-WDR5650 is mounted on each drone to facilitate communication among multiple drones. Each drone is equipped with a light strip, controlled via GPIO on the UP Board 4000® computing board. This setup allows for visualizing the hierarchical reorganization process within the swarm. When a drone’s light strip is illuminated, it indicates that the drone is the current leader of the formation.
@@ V-C Results Analysis
The formation flight trajectories during two instances of hierarchical reorganizations, referred to as Exp. 1 and Exp. 2, are depicted in Fig. 8(a) and 8(b), respectively. Five drones are initially arranged in a regular pentagon with a side length of $1.2m$. Besides, the white connecting lines between each two drones illustrate the current topology of the formation, and the four sub-figures in Fig. 8(a) and 8(b) record the flight trajectory points and topological connection structures during the intervals between two hierarchical reorganizations. In Exp. 1, the formation of hierarchical reorganizations at four nodes when $t=0s,6.4s,11.7s,16.6s,23s$. Similarly, in Exp. 2, the formation also experiences four hierarchical reorganizations, occurring at $t=0s,7s,12.8s,19s,25s$.
TABLE IV: The cost time for hierarchical reorganization of different drone in Exp. 1 and Exp. 2
Drone 0 Drone 1 Drone 2 Drone 3 Drone 4
Exp. 1 ${}^{0}_{1}t(s)$ ${}^{1}_{1}t(s)$ ${}^{2}_{1}t(s)$ ${}^{3}_{1}t(s)$ ${}^{4}_{1}t(s)$
0.037 0.038 0.039 0.038 0.037
${}^{0}_{2}t(s)$ ${}^{1}_{2}t(s)$ ${}^{2}_{2}t(s)$ ${}^{3}_{2}t(s)$ ${}^{4}_{2}t(s)$
0.039 0.037 0.040 0.033 0.038
${}^{0}_{3}t(s)$ ${}^{1}_{3}t(s)$ ${}^{2}_{3}t(s)$ ${}^{3}_{3}t(s)$ ${}^{4}_{3}t(s)$
- - 0.040 - 0.036
Exp. 2 ${}^{0}_{1}t(s)$ ${}^{1}_{1}t(s)$ ${}^{2}_{1}t(s)$ ${}^{3}_{1}t(s)$ ${}^{4}_{1}t(s)$
0.040 0.038 0.039 0.040 0.037
${}^{0}_{2}t(s)$ ${}^{1}_{2}t(s)$ ${}^{2}_{2}t(s)$ ${}^{3}_{2}t(s)$ ${}^{4}_{2}t(s)$
0.037 0.037 $\mathbf{0.047}$ 0.043 -
${}^{0}_{3}t(s)$ ${}^{1}_{3}t(s)$ ${}^{2}_{3}t(s)$ ${}^{3}_{3}t(s)$ ${}^{4}_{3}t(s)$
0.038 - 0.037 0.038 -
TABLE V: The tracking error for different drone in Exp. 1 and Exp. 2
Drone 0 Drone 1 Drone 2 Drone 3 Drone 4
Exp. 1 ${}^{0}e_{x}(m)$ ${}^{1}e_{x}(m)$ ${}^{2}e_{x}(m)$ ${}^{3}e_{x}(m)$ ${}^{4}e_{x}(m)$
0.003 0.004 -0.045 -0.063 -0.038
${}^{0}e_{y}(m)$ ${}^{1}e_{y}(m)$ ${}^{2}e_{y}(m)$ ${}^{3}e_{y}(m)$ ${}^{4}e_{y}(m)$
-0.071 0.051 0.107 0.003 -0.012
Exp. 2 ${}^{0}e_{x}(m)$ ${}^{1}e_{x}(m)$ ${}^{2}e_{x}(m)$ ${}^{3}e_{x}(m)$ ${}^{4}e_{x}(m)$
0.005 0.009 -0.022 -0.027 -0.014
${}^{0}e_{y}(m)$ ${}^{1}e_{y}(m)$ ${}^{2}e_{y}(m)$ ${}^{3}e_{y}(m)$ ${}^{4}e_{y}(m)$
-0.051 0.036 0.044 -0.001 0.027
Figure 9: The trajectory and track error of formation with patterns ’1’.
Figure 10: The trajectory and track error of formation with patterns ’8’.
Figure 11: Hoop-traversing experiment.
Figure 12: The trajectory and track error of Exp. 3.
Table III compares the average flying velocity ${}^{i}\bar{v}$, path length ${}^{i}{path}$ and the number of times each drone served as the leader ${}^{i}{Times}$ in Exp. 1 and 2. These results further validate the outcomes of simulation experiments. During the hierarchical reorganization process, the average velocities and path lengths among the drones are relatively similar, suggesting that the omnidirectional affine formation strategy enhances the in-team coordination of the swarm. The fact that ${}^{i}{Times}$ is non-zero for all drones signifies that each one has acted as the leader at some point, implying that the current leadership evaluation algorithm demonstrates agile responsiveness to collective behaviors such as turning and U-turns. To calculate the cost time during hierarchical reorganizations, the role of each drone at every moment is recorded using $rosbag$ at a frequency of 100Hz. By computing the interpolation between the moments before and after each drone’s role transition, Table IV records the cost time for each drone during hierarchical reorganizations. Data analysis elucidates that the mean durations in Exp. 1 and Exp. 2 are 0.037s and 0.039s, respectively, marginally surpassing the operational control frequency of 30Hz. Notably, the maximum time is 0.047s, which implies that the formation can achieve swift hierarchical reorganization and smooth power transition.
Considering the transitions during hierarchical reorganizations, the actual trajectories and formation center tracking errors for Exp. 1 and 2 are illustrated in Fig. 9(a) and 10(a), respectively. Since a motion capture system is used for positioning during the experimental process, the impact of state estimation accuracy on the actual control effect is disregarded here. As shown in Fig. 9(a) and 10(a), the formation tracking error gradually decreases as it approaches the leader-switching points. Since the tracking performance of individual drones influences the center tracking error, combining this with the distribution patterns of individual drone velocity curves, it can be noted that the center tracking error decreases with a reduction in velocity. Besides, Table V records the tracking errors of different drones in Exp. 1 and 2, and the results show that the tracking errors among different drones are relatively similar. Fig. 9(b) and 10(b) present the expected and actual position curves for the $x$ and $y$ directions of different drones to investigate the distribution pattern of tracking errors. As depicted in Fig. 9(b) and 10(b), the actual position curves consistently lag behind the expected value curves, reflecting the latency inherent in the actual control system.
Figure 13: The velocity curve of the formation in Exp. 3
TABLE VI: The tracking error for different drones and formation transform matr in Exp. 3
Drone 0 Drone 1 Drone 2 Drone 3 Drone 4
Exp. 3 ${}^{0}e_{x}(m)$ ${}^{1}e_{x}(m)$ ${}^{2}e_{x}(m)$ ${}^{3}e_{x}(m)$ ${}^{4}e_{x}(m)$
-0.005 0.008 -0.015 -0.009 -0.007
${}^{0}e_{y}(m)$ ${}^{1}e_{y}(m)$ ${}^{2}e_{y}(m)$ ${}^{3}e_{y}(m)$ ${}^{4}e_{y}(m)$
0.005 0.016 0.023 0.017 0.004
Transform Matrix
$\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\rightarrow\begin{bmatrix}2&0\\ 0.5&0\end{bmatrix}\rightarrow\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\rightarrow\begin{bmatrix}2&0\\ 0.5&0\end{bmatrix}\rightarrow\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$
@@
The hierarchical reorganization mechanism, combined with the affine formation crossing the hoop experiment, is illustrated in Fig 11, referred to as Exp. 3. In Exp. 3, five drones are initially arranged in a regular pentagon with side lengths of $1m$. The formation, relying on the omnidirectional affine formation mechanism, sequentially passed through two red hoops, with the white arrows in the diagram indicating the direction of movement through the hoops. The six sub-figures in Fig. 11 record critical moments of crossing the hoop. At $t=0s$, marking the start of the movement, the formation first moves left and then forward to pass through the left hoop. In order to pass through the red hoop smoothly, the formation transforms into a straight line at $t=5s$, applying affine transformations. It successfully passes through the left red hoop at $t=15s$, returning to the regular pentagonal formation and completing the hierarchical reorganization. Subsequently, the formation reaches the front of the right red hoop at $t=20s$ and transforms into a straight line again at $t=25s$ to pass through the hoop. Finally, the formation returns to the original position at $t=45s$.
TABLE VII: The hierarchical reorganization scenario and formation transform matrix in Exp. 4
@@
The velocity distribution of the formation in Exp. 3 is shown in Fig. 13. In this experiment, the maximum instantaneous velocity of the formation is $1.05m/s$, with the largest velocity difference within the formation being $1.04m/s$ and the smallest being $0.46m/s$. As shown in Fig. 13, compared with Exps. 1 and 2, in the phases of $t=[5-10s,15-20s,25-30s,35-40s]$ in Exp. 3, there are instances where the maximum velocity difference within the formation exceeds the maximum instantaneous velocity. This phenomenon corresponds to the radial formation transformation process, as depicted in Fig. 11. This indicates that during affine transformations, the relative velocities of drones within the formation change, resulting in inter-swarm movements towards each other.
The real drone trajectories and the formation center tracking error in Exp. 3 are shown in Fig. 12(a), which indicates that the formation undergoes dynamic leader switches after crossing through the hoops to complete the turn-around maneuver of the group successfully. Additionally, the tracking error of the formation center trajectory is relatively larger at the beginning of the affine transformation. To analyze the individual drone trajectory tracking error, Table VI records the tracking errors and the formation shape transformation matrix. Compared to Exps. 1 and 2, the individual drone tracking errors in Exp. 3 are relatively more minor, which might be due to the comparatively lower movement speed of the formation in Exp. 3. Furthermore, Fig. 12(b) presents the expected and actual position curves for the $x$ and $y$ directions of different drones in Exp 3. It can be observed that the actual position curves consistently lag behind the expected value curves, consistent with the results of Exps. 1 and 2.
In the experiment involving dense formation changes, as shown in Fig 14, referred to as Exp. 4, five drones are arranged in a regular pentagon with a side length of $1.2m$. The formation undergoes continuous hierarchical reorganizations in the environment depicted in Fig. 7(c). The velocity distribution of the formation in Exp. 4 is shown in Fig. 15, with the maximum instantaneous speed within the formation being $1.92m/s$.
Table VII reflects the hierarchical reorganization scenario and transformation matrices at times labeled in Subfigures of Fig. 14. Combining the information from Fig. 14 and Table VII, it can be observed that at $t=3.1s,14.3s$, the formation changes into a straight line, at $t=0s,13.3s,19.1s$, the formation undergoes scaling, at $t=6.3s,21.2s$, it expands, and at $t=3.1s,6.3s,13.3s$, a dynamic leader switch occurs to respond to rapid maneuvers, such as a formation turning. Exp. 4 illustrates that affine transformations under hierarchical reorganization enable flexible transformations and rapid execution of complex maneuvers, such as formation U-turns.

### v2

@@ IV-B Power-Centric Topology Switching
In hierarchical formation, control under affine transformations typically manifests as leaders and followers employing distinct strategies. The follower’s position, inherently dependent on topological relationships, exhibits a coupling effect between the formation’s inherent topological structures and the temporal difference, affecting the control accuracy. Considering the dynamic change of individual roles within the hierarchical reorganization, invariant topological connections can compromise control accuracy.
Firstly, the impact of formation topology on control precision is analyzed. Eq. (14) presents the classical calculation formula for the follower’s position.
$\dot{\boldsymbol{x}}_{i}=-\frac{1}{\sum_{j\in N_{i}}\omega_{ij}}\sum_{j\in N_{i}}\omega_{ij}(\boldsymbol{x}_{i}-\boldsymbol{x}_{j}-\dot{\boldsymbol{x}}_{j})$ (14)
where $x_{i}$ and $x_{j}$ are the current positions of agent $i$ and $j$, respectively. Due to the presence of errors between the current position and the target position during actual flight, let $\bar{x}_{i}$ and $\bar{x}_{j}$ represent the target positions of i and j, respectively. Then, $\epsilon_{i}=x_{i}-\bar{x}_{i}$ and $\epsilon_{j}=x_{j}-\bar{x}_{j}$. Substituting these into Eq. (1), then
$\sum_{j\in{N_{i}}}\omega_{ij}(\boldsymbol{x}_{i}+\boldsymbol{\epsilon}_{i}-\boldsymbol{x}_{j}-\boldsymbol{\epsilon}_{j})=0$ (15)
Furthermore, the stress formulation can be expressed as:
$\boldsymbol{\epsilon}_{i}-\frac{\sum_{j\in{N_{i}}}\omega_{ij}\boldsymbol{\epsilon}_{j}}{\sum_{j\in{N_{i}}}\omega_{ij}}=-\frac{\sum_{j\in{N_{i}}}\omega_{ij}(\boldsymbol{x}_{i}-\boldsymbol{x}_{j})}{\sum_{j\in{N_{i}}}\omega_{ij}}$ (16)
Based on this, Eq. (14) can be reformulated as:
$(\boldsymbol{\epsilon_{i}}-\frac{\sum_{j\in{N_{i}}}\omega_{ij}{\boldsymbol{\epsilon}_{j}}}{\sum_{j\in{N_{i}}}\omega_{ij}})=-(\boldsymbol{\dot{\epsilon}_{i}}-\frac{\sum_{j\in{N_{i}}}\omega_{ij}\boldsymbol{\dot{\epsilon}_{j}}}{\sum_{j\in{N_{i}}}\omega_{ij}})$ (17)
From Eq. (17), it can be derived that under the follower position calculation equation Eq. (14), the taget position error of agent $i$ is related to the time and the position tracking errors of its neighbors. In response, the following part delves into a power-centric topological switching approach designed to mitigate the influence of topological connections in tandem with the hierarchical reorganization process.
Between the factors of time and neighbors’ position errors, selecting agents with minimal position errors as neighbors can reduce target position calculation errors. The tracking error for leaders is primarily influenced by control errors, whereas for followers, it is affected by both target position calculation errors and control errors. Assuming consistent control performance between leaders and followers, choosing followers as neighbors rather than leaders amplifies the target position calculation error. Furthermore, this calculation error increases with the number of follower neighbors. Based on this, we propose a feasible topology switching strategy to minimize position calculation errors and demonstrate its feasibility.
@@
(Generic Graph Configuration [Lin2016]) For the dynamic formation $\mathcal{G}(\mathcal{V}(t),\mathcal{E}(t))$, if and only if $\mathcal{G}(\mathcal{V}(t),\mathcal{E}(t))$ is $(d+1)-rooted$, the stress matrix $\Omega$ is semi-positive and $rank(\Omega)=n-d-1$.
@@
Theorem IV.3.
(Topology Configuration for Reconfigurable Affine Localizability) Under the Assumption 3, for the dynamic formation $\mathcal{G}(\mathcal{V}(t),\mathcal{E}(t))$, when each follower is topologically connected only to leaders, $\mathcal{G}(\mathcal{V}(t),\mathcal{E}(t))$ can achieve reconfigurable affine localizability.
@@
$\small\boldsymbol{D}_{fl}=\begin{bmatrix}-\mathbf{1}\in\mathbb{R}^{1\times n_{ef}^{1}}&\mathbf{0}\in\mathbb{R}^{1\times n_{ef}^{2}}&...&\mathbf{0}\in\mathbb{R}^{1\times n_{ef}^{n_{f}}}\\ \mathbf{0}\in\mathbb{R}^{1\times n_{ef}^{1}}&-\mathbf{1}\in\mathbb{R}^{1\times n_{ef}^{2}}&...&\mathbf{0}\in\mathbb{R}^{1\times n_{ef}^{n_{f}}}\\ ...&...&...&...\\ \mathbf{0}\in\mathbb{R}^{1\times n_{ef}^{1}}&\mathbf{0}\in\mathbb{R}^{1\times n_{ef}^{2}}&...&-\mathbf{1}\in\mathbb{R}^{1\times n_{ef}^{n_{f}}}\end{bmatrix}$
@@
Denote $\boldsymbol{\bar{r}}=\begin{bmatrix}\boldsymbol{r}_{1}^{T}&\boldsymbol{r}_{2}^{T}&...&\boldsymbol{r}_{n_{l}}^{T}\\ \mathbf{1}&\mathbf{1}&...&\mathbf{1}\end{bmatrix}^{T}$, then $\boldsymbol{\bar{r}}^{T}\boldsymbol{\omega}=0$. However, according to the Lemma 1 in [Zhao2018], $rank(\boldsymbol{\bar{r}})=d+1$, which implies $\boldsymbol{\bar{r}}^{T}$ is of row-full rank. Then, from the properties of vector spaces, only zero vectors exist in the zero space of $\boldsymbol{\bar{r}}^{T}$. According to the definition of $\omega$ in Eq.(2), $\sum_{j=1}^{n_{ef}^{i}}|\omega_{i}^{j}|>0$, which contradicts the above inference obviously. ∎
According to Theorem IV.3, when the hierarchical reorganization is completed through Section IV-A, the topological connection structure in the current configuration is reset, such that each follower establishes a topological connection only with leaders. An example of topological switch is shown in Fig. 3(b).
Figure 4: The system architecture of reconfigurable hierarchical formation.
@@ V-A Simulation
Figure 5: Simulation Process in 3D space.(a) Simulation in Gazebo. (b) Visulization in RViz.
Table I illustrates the impact of hierarchical reorganizations on the different drone path lengths. When employing the formation strategies, Drone 2 exhibits the longest motion path, measuring $121.92m$, while Drone 4 covers the shortest distance at $120.38m$, resulting in a difference of $1.54m$. Conversely, without hierarchical reorganizations, Drone 1 records the longest motion path within the formation at $126.92m$, while Drone 3 covers the shortest distance at $120.88m$, yielding a substantial difference of $6.08m$. This signifies that the proposed strategies can effectively equalize the motion distances of drones, thereby significantly enhancing both the coordination within the formation and the formation’s overall endurance.
@@
With hierarchical reorganization 121.33 121.76 120.38 121.92 121.76
W/O hierarchical reorganization[Zhao2018] 122.28 120.88 126.92 122.60 120.43
@@
With hierarchical reorganization 1.75 1.76 1.73 1.75 1.76
W/O hierarchical reorganization[Zhao2018] 1.56 1.54 1.60 1.56 1.53
Table II presents the influence of hierarchical reorganizations on the average motion speeds of different drones. When employing the proposed strategy, Drone 4 achieves the highest motion speed, recording a speed of $1.76m/s$, while Drone 2 exhibits the slowest motion speed at $1.73m/s$, resulting in a difference of $0.03m/s$. In contrast, without the hierarchical reorganizations, Drone 4 attains the highest motion speed within the cluster at $1.60m/s$, while Drone 2 moves at the slowest speed of $1.53m/s$, resulting in a more substantial difference of $0.07/s$. Furthermore, the overall motion durations for both approaches are $38.93s$ and $37.31s$, respectively. This indicates that, with similar overall motion durations, the proposed strategy significantly reduces the average speed disparities among individual entities, allowing the swarm to maintain similar motion speeds. Consequently, this enhances both the in-team coordination and flexibility of the formation.
@@ V-B Experimental Setup
To verify the performance of the proposed algorithm, we conduct several indoor experiments, as shown in Fig. LABEL:fig:environment. Specifically, Fig. LABEL:fig:environment(a) depicts a scenario set within an obstacle-free dark environment, where five drones execute formation flights in the ${}^{\prime}1^{\prime}$ and ${}^{\prime}8^{\prime}$ patterns. Fig. LABEL:fig:environment(b) and Fig. LABEL:fig:environment(c) present environments featuring a hoop obstacle and a natural obstacle, respectively. In Fig. LABEL:fig:environment(b), the quintet of drones performs a coordinated obstacle avoidance flight through the hoop, whereas in Fig. LABEL:fig:environment(c), they engage in formation flights with free transformations within the natural obstacle environment. The dimensions (length×$\times$×width×$\times$×height) of each drone are $23cm\times 23cm\times 14cm$. The IMU is embedded in each drone’s flight control hardware, referred to as Pixhawk®. Moreover, the NOKOV® motion capture system provides the global position measurements, and an UP Board 4000® computing board running ROS is adopted as the onboard computer. A small TP-LINK® router model TL-WDR5650 is mounted on each drone to facilitate communication among multiple drones. Each drone is equipped with a light strip, controlled via GPIO on the UP Board 4000® computing board. This setup allows for visualizing the hierarchical reorganization process within the swarm. When a drone’s light strip is illuminated, it indicates that the drone is the current leader of the formation.
@@ V-C Results Analysis
TABLE III: The hierarchical reorganization scenario and formation transform matrix in Exp. 4
@@
In the experiment involving dense formation changes, as shown in Fig 6, referred to as Exp. 4, five drones are arranged in a regular pentagon with a side length of $1.2m$. The formation undergoes continuous hierarchical reorganizations in the environment depicted in Fig. LABEL:fig:environment(c). The velocity distribution of the formation in Exp. 4 is shown in Fig. 7, with the maximum instantaneous speed within the formation being $1.92m/s$.
Table III reflects the hierarchical reorganization scenario and transformation matrices at times labeled in Subfigures of Fig. 6. Combining the information from Fig. 6 and Table III, it can be observed that at $t=3.1s,14.3s$, the formation changes into a straight line, at $t=0s,13.3s,19.1s$, the formation undergoes scaling, at $t=6.3s,21.2s$, it expands, and at $t=3.1s,6.3s,13.3s$, a dynamic leader switch occurs to respond to rapid maneuvers, such as a formation turning. Exp. 4 illustrates that affine transformations under hierarchical reorganization enable flexible transformations and rapid execution of complex maneuvers, such as formation U-turns.